OpenStudy (cicilybailey14):

What is the greatest value in the range of y = x2 – 7 for the domain {-2, 0, 1}? Make sure to show all of your work in order to justify your answer and receive credit.

OpenStudy (campbell_st):

substitute each x value to get a corresponding y value. then look at which is the largest... that will be the answer

OpenStudy (cicilybailey14):

Y=-2^2-7=-11 Y=0^2-7=-7 Y=1^2-7=-6 So the first one?